# Symmetry-protection Zeno phase transition in monitored lattice gauge theories

### v1

@@ V Conclusions
In this paper, we have proposed a digital post-selection protocol harnessing passive mid-circuit measurements to both detect and suppress violations of local gauge symmetries in quantum simulations of Abelian LGT models. The error mitigation capacity is not continuous in the monitoring frequency but emerges only after a quantum Zeno transition in the spectrum of an associated Liouvillian superoperator, which separates the protected phase from a regime in which the measurements induce a dynamics that quickly thermalizes to an infinite temperature ensemble. This framework based on the quantum Zeno effect unifies many different proposals that exploit stochastic dynamics to protect the gauge symmetries in both digital and analog quantum simulation protocols, such as random gauge transformations [31], engineered dissipation [30], and syndrome measurements [36]. Crucially, despite the underlying Liouvillian being identical, each protocol results in a different unraveling in quantum trajectories. We exemplified this feature by comparing the stroboscopic projective measurements in a digital quantum simulation (the DPS protocol), with continuous weak measurements of the local charges in a continuous-time setup. The ensemble averages of the two stochastic processes are the same while their distributions are surprisingly different.
@@
Moreover, we showed that a simple error-correcting scheme harnessing the natural redundancy in LGTs [37] can be easily integrated with DPS circuits. Instead of using a qubit redundancy, we rely purely upon the constraint imposed by the local gauge symmetries to detect and correct the incoherent bit-flip errors with a feedback mechanism, while at the same time suppressing a coherent gauge drift. Though weaker than similar fully error-correcting schemes [37], it is also significantly less demanding as it only requires at most $N$ ancilla qubits for the measurements, instead of auxiliary computational qubits and the associated heavier logical gates. Thus, our approach will be useful during the transition between NISQ and fault-tolerant devices, where the number of qubits and their connectivity are still limiting factors.
In our analysis, we focused on the ${\mathbb{Z}}_{2}$ gauge group in 1+1 dimensions. The extension to larger dimensions is conceptually trivial, even though the measurements of the local charges involve a larger number of entangling gates to map their parity to the ancillary qubits [36]. When considering different LGTs, a popular choice is a truncated $U(1)$ gauge group such as the $S=1/2$ quantum link model [6, 12, 34]. As it shares the same Hilbert space of the ${\mathbb{Z}}_{2}$ LGT, we expect similar results to those we presented in this paper, with minor modifications to accommodate for the different forms of the local symmetries. A more interesting question is how to adapt the construction to larger groups, possibly non-Abelian. In this situation, a natural platform choice is represented by qudits, where a single quantum register encodes either the full symmetry group of a link or an appropriate truncation [84, 85, 15]. As the local charges adopt more values, the ancillary registers must also be encoded in qudits with a sufficient number of levels. These features complicate the design of the DPS circuit but do not change its core idea, although, for non-Abelian groups, results may depend on the order in which the local symmetry operators are measured.
@@
An aspect that we did not consider in this paper is the effect of a dephasing error channel. Indeed, we showed that DPS with error correction can suppress both coherent and bit-flip gauge-breaking errors, but it is insensitive to dephasing as such an error does not affect Gauss’ laws. A possible approach to refine our proposal is to use qubit redundancy with majority votes to correct for dephasing errors and the local charge measurements for bit-flips [37], at the price of implementing more costly error-corrected logical gates. Moreover, in our analysis, we assumed that the error accumulated in each layer of the circuit is proportional to the discretization step $\Delta t$. Although this is irrelevant in the DPS scheme, it affects the performance dependence on the number of Trotter steps and is necessary for the connection with the continuous-time evolution approach. Typically, we need to consider the error rates of nonparametric gates, such as CNOT, which is independent of the time step. In particular, this will appear also in the circuit that embeds the parity of the local charges in the ancilla qubits, resulting in an imperfect error suppression even when the correct parity is measured. All these effects can be incorporated into a more detailed analysis of DPS with a hardware-specific noise model and elementary gates.
@@ Acknowledgements
We thank J. Mildenberger, L. Spagnoli, A. Russomanno, and M. Burrello for fruitful discussions. This project has received funding from the European Union’s Horizon Europe research and innovation programme under grant agreement No 101080086 NeQST and from European Union - NextGeneration EU, within PRIN 2022, PNRR M4C2, Project TANQU 2022FLSPAJ [CUP B53D23005130006]. P.H. has further received funding from the Italian Ministry of University and Research (MUR) through the FARE grant for the project DAVNE (Grant R20PEX7Y3A), from the Swiss State Secretariat for Education, Research and lnnovation (SERI) under contract number UeMO19-5.1 , from the QuantERA II Programme through the European Union’s Horizon 2020 research and innovation programme under Grant Agreement No 101017733, from the European Union under NextGenerationEU, PRIN 2022 Prot. n. 2022ATM8FY (CUP: E53D23002240006), from the European Union under NextGenerationEU via the ICSC – Centro Nazionale di Ricerca in HPC, Big Data and Quantum Computing. This project has been supported by the Provincia Autonoma di Trento and Q@TN, the joint lab between the University of Trento, FBK—Fondazione Bruno Kessler, INFN—National Institute for Nuclear Physics, and CNR—National Research Council. Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Commission. Neither the European Union nor the granting authority can be held responsible for them. AB would like to thank the Institut Henri Poincaré (UAR 839 CNRS-Sorbonne Université) and the LabEx CARMIN (ANR-10-LABX-59-01) for their support.
@@
Appendix B Trajectory fluctuations in analog quantum simulations
@@
Appendix C Trajectory post selection
In Sec. IV, we adopted an outlier exclusion method to automatically filter out trajectories in which the error correction failed because of non-detectable errors affecting neighboring symmetry generators. Here we give more details on the method used for the data in in Fig. 6 (dashed green line and triangular markers). At each Trotter step, we check the distance between an individual trajectory and the average in units of the standard deviation of the observable. If at any point this distance is larger than a certain threshold, we discard the trajectory, assuming that a non-detectable error occurred. This process can be iterated several times to further refine the ensemble of trajectories used to compute the corrected dynamics.

### v3

@@ VI Conclusions
In this paper, we have proposed a dynamical post-selection protocol harnessing passive mid-circuit measurements to both suppress (coherent) and correct (incoherent) violations of local gauge symmetries in quantum simulations of Abelian LGT models. The error mitigation capacity is not continuous in the monitoring frequency but emerges only after a quantum Zeno transition in the spectrum of an associated Liouvillian superoperator, which separates the protected phase from a regime in which the measurements induce a dynamics that quickly thermalizes to an infinite temperature ensemble. This framework based on the quantum Zeno effect unifies many different proposals that exploit stochastic dynamics to protect the gauge symmetries in both digital and analog quantum simulation protocols, such as random gauge transformations [32], engineered dissipation [31], and syndrome measurements [38]. Crucially, despite the underlying Liouvillian being identical, each protocol results in a different unraveling in quantum trajectories. We exemplified this feature by comparing the stroboscopic projective measurements in a digital quantum simulation (the DPS protocol), with continuous weak measurements of the local charges in a continuous-time setup. The ensemble averages of the two stochastic processes are the same while their distributions are surprisingly different. Importantly, noisy measurements do not spoil the gauge protection, although the Zeno-like regime does not hold any longer to arbitrary large monitoring frequency.
@@
Moreover, we showed that a simple error-correcting scheme harnessing the natural redundancy in LGTs [39] can be easily integrated with DPS circuits. It relies purely upon the constraint imposed by the local gauge symmetries to detect and correct the incoherent bit-flip errors with a feedback mechanism. Though weaker than similar fully error-correcting schemes [39, 41], it is also significantly less demanding as it only requires at most $N$ ancilla qubits for the measurements, instead of auxiliary computational qubits and the associated heavier logical gates. It showcases how coherent and incoherent errors are handled differently within the same quantum circuit: first, frequent measurements suppress the coherent gauge drift, and, second, the active feedback builds upon the measurement results to compensate for incoherent spin-flips. Hence, there is no need to implement further gauge protection schemes to suppress coherent errors, typically also requiring entangling gates, if mid-circuit measurements are sufficiently reliable. Considering that the exponential overhead of error mitigation techniques limits their scalability in NISQ devices, it is of fundamental importance to understand how far we can push the present generation of quantum technologies in the quest for fault tolerance. Our approach provides valuable insights into the transition between NISQ and fault-tolerant devices, where quantum resources are still limiting factors and error-mitigation/correction strategies also need to cope with their possibly undesired side effects on quantum simulations. It stands as an important open research question whether error-mitigated NISQ devices can provide a real advantage for practical problems or give a precious contribution to guide the development towards more advanced scalable algorithms.
In our analysis, we focused on the ${\mathbb{Z}}_{2}$ gauge group in 1+1 dimensions to illustrate our scheme in the simplest scenario. The extension to larger spacetime dimensions is conceptually trivial, even though the measurements of the local charges involve a larger number of entangling gates to map their parity to the ancillary qubits [38, 88]. To consider different abelian LGTs with larger groups, we studied the effective Liouvillian spectrum associated with the Zeno transition in models in one spatial dimension with $\mathbb{Z}_{3}$ and $U(1)$ symmetries, the latter truncated to a quantum link model with $S=1$ [89, 90, 6, 12, 35]. As shown in Appendix B, we obtained qualitatively similar results with ${\mathbb{Z}}_{2}$ LGT, even considering noisy measurements.
When extending our protocol to larger groups, the main challenge is to design efficient measurement and correction circuits. Errors are still detected by monitoring the local charges [38] but inverting the errors becomes more complicated because of the richer structure of possible gauge-breaking terms. In this situation, a natural platform choice is represented by qudits, where a single quantum register encodes either the full symmetry group of a link or an appropriate truncation [91, 92, 15]. As the local charges adopt more values, the ancillary registers must also be encoded in qudits with a sufficient number of levels [88]. These features complicate the design of the DPS circuit but do not change its core idea, although, for non-Abelian groups, results may depend on the order in which the local symmetry operators are measured. On a positive note, using qudits would reduce the number of entangling gates needed for the symmetry verification, with respect to qubit devices, thus making the impact of measurement-induced noise less severe. Larger groups also impose stricter constraints on the physical states, meaning that monitoring the local charges would suppress a broader range of errors.
@@
An aspect that we did not consider in this paper is the effect of a dephasing error channel. Indeed, we showed that DPS with error correction can suppress both coherent and bit-flip gauge-breaking errors, but it is insensitive to errors not affecting Gauss’ laws. A possible approach to refine our proposal is to use qubit redundancy with majority votes to correct dephasing errors and the local charge measurements for bit-flips [39, 41], at the price of implementing more costly error-corrected logical gates. Moreover, in our analysis, we assumed that the error accumulated in each layer of the circuit is proportional to the discretization step $\Delta t$. Although this is irrelevant in the DPS scheme, it will likely affect its performance dependence on the number of Trotter steps. When considering the error rates of nonparametric gates, such as CNOT, the Zeno protection will appear only in an optimal window of measurement frequency ranges. This is essentially what emerges when the circuit coupling the ancilla with the computational qubits is noisy, as we have shown in Sec. IV, resulting in an imperfect but still effective error suppression. In the future, it will be interesting to perform a thorough analysis with realistic error and noise sources to fully assess the efficiency of DPS in practical scenarios.
@@ Acknowledgements
We thank J. Mildenberger, L. Spagnoli, A. Russomanno, and M. Burrello for fruitful discussions. This project has received funding from the European Union’s Horizon Europe research and innovation programme under grant agreement No 101080086 NeQST and from European Union - NextGeneration EU, within PRIN 2022, PNRR M4C2, Project TANQU 2022FLSPAJ [CUP B53D23005130006]. P.H. has further received funding from the Italian Ministry of University and Research (MUR) through the FARE grant for the project DAVNE (Grant R20PEX7Y3A), from the Swiss State Secretariat for Education, Research and Innovation (SERI) under contract number UeMO19-5.1 , from the QuantERA II Programme through the European Union’s Horizon 2020 research and innovation programme under Grant Agreement No 101017733, from the European Union under NextGenerationEU, PRIN 2022 Prot. n. 2022ATM8FY (CUP: E53D23002240006), from the European Union under NextGenerationEU via the ICSC – Centro Nazionale di Ricerca in HPC, Big Data and Quantum Computing. This project has been supported by the Provincia Autonoma di Trento and Q@TN, the joint lab between the University of Trento, FBK—Fondazione Bruno Kessler, INFN—National Institute for Nuclear Physics, and CNR—National Research Council. Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Commission. Neither the European Union nor the granting authority can be held responsible for them. AB would like to thank the Institut Henri Poincaré (UAR 839 CNRS-Sorbonne Université) and the LabEx CARMIN (ANR-10-LABX-59-01) for their support.
Data availability statement
All data supporting this work can be found in [95].
@@
Appendix B Effective Liouvillian spectrum in $\mathbb{Z}_{3}$ and $U(1)$ gauge groups.
Our analysis is not restricted to the simple ${\mathbb{Z}}_{2}$ model discussed in the main text. In this section, we extend the investigation of the Liouvillian spectrum associated with the continuous-time limit of the DPS protection protocol to LGTs in one spatial dimension with $\mathbb{Z}_{3}$ and $U(1)$ symmetries. For both, the structure of the Hamiltonian is the same as Eq. (1) but the operators acting on the gauge degrees of freedom and the local charges need to be changed to take into account the larger groups.
B.0.1 $\mathbb{Z}_{3}$ group
A LGT with discrete $\mathbb{Z}_{N}$ gauge group is a generalization of the simpler model analyzed in the main text. We consider generalized Pauli matrices (clock operators) satisfying the algebra
$Q^{N}=P^{N}=\mathbb{1}\ ,$ (24a)
$QP=PQ\textrm{e}^{i\frac{2\pi}{N}}\ ,$ (24b)
$Q|m\rangle=\textrm{e}^{i\frac{2\pi}{N}m}|m\rangle\ .$ (24c)
Here, we focus on $N=3$. After a Jordan–Wigner transformation on the fermions (matter sites), the Hamiltonian (1) then becomes
$H_{\mathbb{Z}_{3}}=$ $J\sum_{n}\left(\sigma^{+}_{n}P_{n,n+1}\sigma^{-}_{n+1}+{\rm h.c}\right)$ (25)
$+$ $\frac{f}{2}\sum_{n}(Q_{n,n+1}+Q^{\dagger}_{n,n+1})+\frac{\mu}{2}\sum_{n}(-1)^{n}\sigma^{z}_{n}\ .$
The gauge transformations are [5]
$G_{n}=Q_{n-1,n}Q^{\dagger}_{n,n+1}\textrm{e}^{-i\frac{\pi}{3}(1+\sigma^{z}_{n})}$ (26)
and their eigenvalues are the cubic roots of unity $g_{n}=\textrm{e}^{i\frac{2\pi}{3}m}$, with $m=0,1,2$.
B.0.2 Truncated $U(1)$
$U(1)$ is the gauge group of quantum electrodynamics [90], making it one of the most popular targets for quantum simulations of LGT. Being a continuous group, we need to truncate it to map it on the discrete degrees of freedom of quantum hardware. Here, we adopt the quantum-link-model formulation for the gauge fields [90, 89]. In practice, we consider the operators $S^{z}$ and $U$ acting on the links, such that
$S^{z}|m\rangle=m|m\rangle\ {\rm with}\ m=-S,-S+1,\dots,S\ ,$ (27)
and
$U|m\rangle=|m+1\rangle\ {\rm if}\ m<S\ .$ (28)
Using Kogut–Susskind staggered fermions, followed by a Jordan–Wigner transformation to hard-core bosons, the Hamiltonian of the theory is
$H_{U(1)}=$ $J\sum_{n}\left(\sigma^{+}_{n}U_{n,n+1}\sigma^{-}_{n+1}+{\rm h.c}\right)$ (29)
$+$ $f\sum_{n}\left(S^{z}_{n,n+1}\right)^{2}+\frac{\mu}{2}\sum_{n}(-1)^{n}\sigma^{z}_{n}\ ,$
while the local Gauss’ law generators read
$G_{n}=S^{z}_{n,n+1}-S^{z}_{n-1,n}-\frac{1}{2}((-1)^{n}+\sigma^{z}_{n})\ .$ (30)
The physical subspace is the one with $G_{n}=0$ and the $(-1)^{n}$ factor takes into account the different values of the charge on even and odd sites. We truncate the gauge fields to $S=1$, meaning that the electric field can take three different values on each link. The Hilbert space is, therefore, the same as the $\mathbb{Z}_{3}$ LGT but the local symmetries and the constraints they impose are different.
Figure 10: (a-d): Liouvillian spectrum as a function of the measurement frequency in the presence of unitary errors for the $\mathbb{Z}_{3}$ (a-b) and $U(1)$ (c-d) gauge groups. The branching in the spectrum signals the onset of the quantum Zeno phase that protects gauge invariance and survives also for finite measurement fidelity, see panels (b) and (d). The opacity of the markers reflects the density of states. (e-g): example of dynamics in the deconfined phase of a 1+1d $\mathbb{Z}_{3}$ LGT model. The Zeno protection with $\gamma=20\lambda$ (f) recovers the target dynamics (e), while it is almost completely lost (g) without it. Simulation parameters: $f=0.5J$, $\mu=0.1J$, $\lambda=0.2J$. In panels (a-d) the chain consists of $N=3$ matter sites, while in panels (e-f) $N=7$.
B.0.3 Liouvillian spectrum
Following the gauge-symmetry-violating Hamiltonian used in the main text, here we consider two sources for coherent errors, rotations of the gauge fields and unassisted hopping between matter sites. The latter is independent of the gauge group and is identical to the one in the main text. On-site electric field rotations take the form $(P_{n,n+1}+P^{\dagger}_{n,n+1})$ for $\mathbb{Z}_{3}$ and $(U_{n,n+1}+U^{\dagger}_{n,n+1})$ for $U(1)$.
To investigate the presence of a quantum Zeno transition, we consider an effective Lindbladian where the unitary evolution is given by the model Hamiltonian and the coherent error terms with a strength $\lambda$, while the collapse operators correspond to the local charges in Eqs. (26) and (30). Measuring the generators of gauge transformations for larger groups is technically more complicated than the simple stabilizer of the ${\mathbb{Z}}_{2}$ group. A circuit implementation can be found, for instance, in Ref. [38]. We also consider the situation in which the measurements of the local charges have finite fidelity, by adding further on-site dissipative noise with a rate $\gamma(1-\mathcal{F})$, where $\gamma$ is the measurement frequency and $\mathcal{F}$ their average fidelity.
Figure 10 summarizes our results: the appearance of the quantum Zeno phase in the Liouvillian spectrum, see panels (a-d) is qualitatively similar to what we observed in Sec. III for both $\mathbb{Z}_{3}$ and $U(1)$ groups, even though the critical measurement frequency slightly changes. The Zeno phase survives also with finite measurement fidelity (here $\mathcal{F}=0.99$) on a finite range of $\gamma$, as observed for the ${\mathbb{Z}}_{2}$ LGT model. When looking at a specific example of gauge-invariant dynamics one wishes to recover, the effect of the measurements is again similar to what we observed for the ${\mathbb{Z}}_{2}$ model. We report an instance for the $\mathbb{Z}_{3}$ group in panels (e-g) of Fig. 10, where a dramatic improvement in the Zeno phase becomes clear with respect to the bare evolution subject to coherent errors.
Appendix C Trajectory fluctuations in analog quantum simulations
@@
Appendix D Trajectory post-processing
In Sec. V, we adopted an outlier exclusion method to automatically filter out trajectories in which the error correction failed because of non-detectable errors affecting neighboring symmetry generators. Here we give more details on the method used for the data in in Fig. 8 (dashed green line and triangular markers). At each Trotter step, we check the distance between an individual trajectory and the average in units of the standard deviation of the observable. If at any point this distance is larger than a certain threshold, we discard the trajectory, assuming that a non-detectable error occurred. This process can be iterated several times to further refine the ensemble of trajectories used to compute the corrected dynamics.